OpenStudy (anonymous):

Is there an easy way to determine whether to use Permutations, Combinations or some differentiation thereof? For Example, A coach randomly selects 3 swimmers from a team of 8 to swim in a heat. What is the probability that she will pick the three strongest swimmers?

OpenStudy (lgbasallote):

this is combinatorics? or whatever that thing was called

OpenStudy (anonymous):

Theoretical and Experimental probability.

OpenStudy (lgbasallote):

hmm @zepp wanna help?

OpenStudy (lgbasallote):

i called you coz you rhyme with help lol

OpenStudy (zepp):

Oh my goodness, @lgbasallote You know that I hate probabilities :(

OpenStudy (lgbasallote):

oh wow..i didnt know that..what are the odds! lol

OpenStudy (unklerhaukus):

Permutations are used when the order is important combinations are used the ordering is not significant

OpenStudy (anonymous):

what unklerhaukus said. you really have to tell from the context. and of course sometimes it is neither, or a combination of the two

OpenStudy (anonymous):

of course i could plug this into my calculator using the nCr key but how would i solve this by hand?

OpenStudy (anonymous):

hold the phone i didn't read the question, but now i see you have to pick exactly the three strongest swimmers out of the 8

OpenStudy (anonymous):

there are two ways to do this. one way is to compute how many ways she can pick 3 out of 8, and that is \(\dbinom{8}{3}=\frac{8\times 7\times 6}{3\times 2}=8\times 7=54\) that will be the denominator then for the numerator you take the number of ways to pick the three strongest out of the 8 there is only one way to do that your answer using this method is \[\frac{1}{54}\]

OpenStudy (anonymous):

56 haha

OpenStudy (anonymous):

8x7=56 but thanks i think i have it now

OpenStudy (anonymous):

the other way is to compute as follows probability she picks one of the strongest as first pick is \(\frac{3}{8}\) then probilility the next pick is from the strongest 3 is \(\frac{2}{7}\) and finally the last pick has probability \(\frac{1}{6}\) multiply these together and you get \[\frac{3}{8}\times \frac{2}{7}\times \frac{1}{6}=\frac{1}{\text{whatever eight times seven}}\] is

OpenStudy (anonymous):

either method works, and it requires not thinking "do i use permutations or combinations" but rather understanding what exactly the question is asking also helps to know \(8\times 7\)

OpenStudy (anonymous):

Alright man thanks a bunch but.... un mas?

OpenStudy (anonymous):

post

OpenStudy (anonymous):

There are 7 books numbered 1-7 on the summer reading list. Peter randomly chooses 2 books. What is the probability that peter chooses books numbered 1 and 2?

OpenStudy (anonymous):

similar to last one

OpenStudy (anonymous):

picks one or two on first choice with probability \(\frac{2}{7}\) then picks the other (given that fist pick was either one or two) with probability \(\frac{1}{6}\) multiply them together because this is an "and" statement and get \[\frac{2}{7}\times \frac{1}{6}\] whatever that is

OpenStudy (anonymous):

other method you have \(\dbinom{7}{2}=\frac{7\times 6}{2}\) ways to pick any two from a set of 7, and only one way to pick both one and two in two picks answer is \[\frac{1}{21}\] for either method

OpenStudy (anonymous):

ah i see, but how does the and statement change it, what if it was "or" instead of "and"?

OpenStudy (anonymous):

depends on the context, but sometimes "or" means "add" or rather add and subtract the intersection it is hard to answer a general question like that because english language is not math, so it is ambiguous